OpenStudy (anonymous):

Find the slope-intercept equation of the line that has the given characteristics. Slope -7 and y-intercept (0,4) The equation is y=

OpenStudy (anonymous):

apply the formula\[y-y1=m(x-x1)\]

OpenStudy (anonymous):

where -7 is slope and y1=4 and x=0

OpenStudy (anonymous):

y - y1 = m (x-x1) y - 4 = -7(x - 0) y - 4 = -7x y = -7x + 4